OpenStudy (anonymous):

What is the product of 7the square root of 10. • 3the square root of 5.? Simplify if possible. @Nurali

OpenStudy (nurali):

\[7\sqrt{10}*3\sqrt{5}\]

OpenStudy (anonymous):

yes

OpenStudy (nurali):

\[7*3*5\sqrt{2}\] \[105\sqrt{2}\]